11x- 8y=-48
Writing Linear 3quation's

Write the Slope intercept form of the equation of Each line
in PLz have all the Step's Iam Very Slow in Math :P

Answer by mananth(16949)

11x- 8y=-48
you have to rewrite the equation in slope intercept form y = mx +b where m is the slope and b = y intercept.
11x-8y=-48
add -11x to both sides
11x-11x-8y= -11x-48
-8y=-11x-48
divide both sides by -8
-8y/-8 = -11x/-8 -48/-8
y=(11/8)*x +8
..
compare this with the equation y = mx +b
m = (11/8) = slope
b = 8 the y intercept.
..
m.ananth@hotmail.ca